# The Influence of V-Defects, Leakage, and Random Alloy Fluctuations on the Carrier Transport in Red InGaN MQW LEDs

### v1

@@ I Introduction
Figure 1: (a) Structure of red LED from Ref [Zhang et al. (2020)]. (b) Cross-section of a V-defect and threading dislocation line at the of the V-defect. The TD dislocation line is modeled as a 30 nm diameter cylinder with a high density of fast nonradiative traps, and a short nonradiative lifetime value in this region. (c) Conduction band potential plotted in log scale. The strong polarization-related and band offset barriers for the c-plane QWs result in carriers flowing through semipolar planes with much-reduced energy variations. (d) Computed band extrema across the vertical LED structure away from the V-defect, taking alloy fluctuations into account (green and violet) or not (blue and red) curves, respectively. In the case of alloy fluctuations, band potential is spatially-averaged over the QW in-plane area. Random fluctuations aid carrier flow by decreasing the effective potential of quantum barriers Li et al. (2017).
@@ II Methodology
For this research, we employed an in-house two-dimensional drift-diffusion charge-control (2D-DDCC) solver to model carrier transport and the electrical properties of $\mu$-LEDs (Fig. 2) Wu et al. (2012). This solver is particularly well-suited for analyzing V-defects and random alloy fluctuations, both of which are crucial for understanding device behavior. First, the Poisson equation (1) was used to calculate the electrostatic potential ($\psi$) based on the charge distribution. Charge sources include free electrons ($n$) and holes ($p$), as well as ionized acceptors and donors, $N_{A}^{-}$ and $N_{D}^{+}$, respectively, while $\rho_{\text{pol}}$ accounts for the polarization charge density, which depends on the local indium composition. Furthermore, the solver integrates the Localization Landscape (LL) theory Li et al. (2017); Filoche et al. (2017); Piccardo et al. (2017), an advanced method that accounts for both quantum confinement effects and alloy disorder in determining the effective quantum potential $\frac{1}{\tau_{e,h}}$. This is achieved by substituting (2) for the conventional Schrödinger equation Filoche et al. (2017). The model integrates in the energy space using the Fermi-Dirac distribution ($f_{n},f_{p}$) and the local density of states for electrons ($N_{c}$) and holes ($N_{v}$) in Eqs. (3) and (4) to compute carrier concentrations, where $E_{fn}$ and $E_{fp}$ denote the quasi-Fermi levels for electrons and holes, respectively.
@@
To simulate random alloy fluctuations, we used a random number generator combined with Gaussian averaging methods to distribute indium (In) and gallium (Ga) atoms across the lattice sites within the multiple quantum wells (MQWs). This arrangement effectively simulates local compositional variations and their impact on essential metrics such as bandgap energy and polarization, ensuring accurate device characterization. Our model also accounts for the impact of tail states, which are carrier states below the bandgap that arise from crystal structure imperfections, strain-induced bandgap changes, or impurities. These tail-state effects have been widely discussed in polycrystalline, doped, or amorphous semiconductor materials. Such states provide additional carrier percolation paths, significantly influencing carrier dynamics. The origin of tail states in this paper will be discussed later. To incorporate tail states, we define the density of tail states, $N_{tail}(E)$, represented by a Gaussian distribution [Eq. (12)], as shown in Fig. 3(d). This equation describes the peak density of tail states $N_{tail}$, the energy center $E_{tail}$, and the standard deviation $\sigma_{tail}$.
@@
Here, $N_{(c,v),total}(E)$ replaces $N_{(c,v)}$ in Eqs. (3) and (4), allowing carriers to exist in these tail states and hop through them inside the energy bandgap, as illustrated in Fig. 3(d). Tail states have long been studied in disordered or doped semiconductors. Specifically, tail states have been identified and computed in doped GaAs, where they originate from the local electric fields of ionized impurities Casey Jr and Stern (1976). These tail-state models have been successfully implemented in our software to describe carrier transport in organic and classical semiconductors Kirchartz et al. (2011); Kung et al. (2017); Huang et al. (2020, 2023). Similar functions are now applied here to describe carrier transport in tail states near the TD region.
@@ II.1 Parameter setting
Table 3: The parameters for tail states density and FWHM of Gaussian distribution in threading dislocation lines and V-defect sidewalls. $E_{tail}$ is zero in the setting. The value used here gives a satisfactory fit to the WPE curve shown in Ref. [Zhang et al. (2020)].
@@
Further material parameters, such as bandgap energies, doping concentrations, and recombination coefficients, are provided in Table 1. The Auger-Meitner coefficient $C_{0}$ is chosen to be significantly larger in the InGaN QW region, as it is well known that alloy disorder substantially increases this coefficient compared to pure GaN Shahmohammadi et al. (2017); Ryu et al. (2020). This choice of $C_{0}$ also provides better agreement with the experimental results discussed later.
@@
The effect of varying indium compositions in the V-defect sidewall QWs on red InGaN LED performance was evaluated. The device structure maintains a constant indium composition of 29% in the c-plane QWs. The thin (0.67 nm) V-defect sidewall QWs were designed with an indium composition of 17%, along with 3.62 nm thick sidewall QBs. These variations lead to corresponding changes in the bandgap energies and polarization properties of the sidewall QWs, which can impact carrier dynamics and recombination efficiency. In the TD center, within 30 nm regions, there is a high density of deep-level traps and shallow-level tail states. The deep-level trap state parameters are listed in Table 2, while the tail state parameters are shown in Table 3.
@@
Our findings indicate that tail states significantly influence carrier transport Kung et al. (2017); Huang et al. (2020), particularly when strong polarization-induced fields and large band offsets hinder vertical carrier injection into the QWs, as in red LEDs. This disrupts the ideal carrier flow over barriers, as illustrated by the band structure model in Fig. 3(c). The altered energy landscape redirects carriers, increasing leakage to non-radiative pathways and promoting recombination at NR states near dislocations. Furthermore, the inclusion of tail states (Table 3) modifies the density of states (DOS), as shown in Fig. 3(d). The additional energy levels introduced by tail states increase the probability of carrier trapping, enhancing nonradiative recombination in dislocation regions and reducing radiative efficiency. These effects on the DOS highlight the critical role of tail states in shaping the electronic properties and overall optical performance of the LED.
Figure 3: Simulation models: (a) Compositional map of the combined model incorporating alloy fluctuations and V-defects. Random alloy fluctuations are illustrated in the inset. Indium composition is 17% at V-defect sidewalls QWs, averaged to 29% at c-plane MQWs, and (b) tail states along TD centers and sidewalls, acting as transport and recombination pathways, reducing the turn-on voltage and radiative efficiency, inducing a leakage current below turn-on. (c) Band structure showing carrier transport through tail states at low current density, leading to an improved carrier injection at low voltage, which is identified as carrier leakage. (d) Position of tail states in the density of states across the bandgap.
@@ III.1 Carrier Injection Mechanism Through V-defects
In Fig. 5(a-b), V-defects and random alloy fluctuations are incorporated into the third model. Since the sidewalls of the QWs are semipolar planes with minimal piezoelectric polarization, the V-defects provide an additional channel for carrier injection. As a result, carrier flow into the c-plane QWs is more efficient through the sidewalls of the V-defects. The reduced barriers, caused by the lower piezoelectric polarization, allow for a more concentrated current density at the V-defect sites, where the voltage at a current density of 20 A/cm2 significantly drops to 2.85V. Additionally, carriers trapped along the TD lines may lead to nonradiative recombination losses. Although the forward voltage decreases significantly, it remains higher than the experimental value of approximately 2.1 V at 1 A/cm2. This discrepancy arises because, although the V-defect $\{10\overline{1}1\}$ semipolar plane ($\sim 62^{\circ}$) exhibits nearly zero polarization at the interface, it typically features a thinner QW with a lower indium composition, similar to that of blue/violet MQWs. The theoretical turn-on voltage of blue LEDs is approximately 2.72V ($\hbar\omega=2.72eV$). Consequently, the sidewall QW aids carrier injection once it becomes conductive. Even if the sidewall QW had a higher indium composition (22%) instead of the 17% used, the voltage would decrease slightly but still would not reach the low turn-on voltage due to the increased band offset in the QW and the higher series resistance resulting from the presence of nine deeper sidewall QWs. Therefore, incorporating a tail state current transport model is necessary.
Figure 5: Computations of current density distributions of (a, c, e) electrons and (b, d, f) holes at applied voltages of V = 2.85V and 2.66V for the third model (compositional fluctuations + V-defects, (a, c) and (b, d), and for the fourth model (addition of tail states (e, f)). V-defects improve carrier flow while tail states increase leakage and nonradiative recombination in the defect region. However, holes are then leaking to lower QWs so that more QWs will emit light. Also, a lower $V_{for}$ is needed to reach a current density of 20 A/cm2 compared to the model without tail states.
@@
When the voltage is lowered to 2.66V, fewer carriers are injected in the third model due to the absence of defect-mediated pathways, requiring a higher voltage to achieve comparable injection levels (Fig. 5(c, d)). However, in the fourth model, which includes tail states, the current density reaches 20 A/cm2 at this voltage (Fig. 5(e, f)). These tail states are incorporated into the V-defect region in the fourth model, with $N_{tail}$ being higher at the defect center (the location of the dislocation line, Fig. 3(b)). The higher current densities in the TD area (Figs. 5(c) and 5(f)) indicate that these tail states increase the current, particularly at lower voltages, while also leading to greater nonradiative recombination, as the TD area contains fast nonradiative recombination traps. The tail states also create leakage paths toward nonradiative centers, reducing overall efficiency. However, they facilitate carrier injection into the QWs at low voltages, as some leakage current diffuses into the QWs. Once carriers begin injecting into the c-plane quantum well (QW), they screen the polarization field, particularly at the corners of the c-plane/semipolar plane interface. This screening effect reduces the injection barrier in the c-plane region. Comparing Figs. 5(d) and 5(f) for hole injection at the same voltage, we observe slightly stronger vertical injection in the c-plane or semipolar plane, especially at the QW corners. Additionally, the J-V curves in Figs. 7 and 8 indicate that this change is not limited to leakage current below the turn-on voltage: the entire J-V curve shifts to a lower voltage.

### v2

@@ I Introduction
Figure 1: (a) Bandgap distribution map of the simulated 9QW Red LED structure with random alloy fluctuations and V-defects. (b) Cross-section of a V-defect and threading dislocation line at the of the V-defect. The TD dislocation line is modeled as a 30 nm diameter cylinder with a high density of fast nonradiative traps, and a short nonradiative lifetime value in this region. (c) Conduction band potential plotted in log scale. The strong polarization-related and band offset barriers for the c-plane QWs result in carriers flowing through semipolar planes with much-reduced energy variations. (d) Computed band extrema across the vertical LED structure away from the V-defect, taking alloy fluctuations into account (green and violet) or not (blue and red) curves, respectively. In the case of alloy fluctuations, band potential is spatially-averaged over the QW in-plane area. Random fluctuations aid carrier flow by decreasing the effective potential of quantum barriers Li et al. (2017).
@@ II Methodology
For this research, we employed an in-house two-dimensional drift-diffusion charge-control (2D-DDCC) solver to model carrier transport and the electrical properties of $\mu$-LEDs (Fig. 2) Wu et al. (2012). This solver is particularly well-suited for analyzing V-defects and random alloy fluctuations, both of which are crucial for understanding device behavior. First, the Poisson equation (1) was used to calculate the electrostatic potential ($\psi$) based on the charge distribution. Charge sources include free electrons ($n$) and holes ($p$), as well as ionized acceptors and donors, $N_{A}^{-}$ and $N_{D}^{+}$, respectively, while $\rho_{\text{pol}}$ accounts for the polarization charge density, which depends on the local indium composition. Furthermore, the solver integrates the Localization Landscape (LL) theory Li et al. (2017); Filoche et al. (2017); Piccardo et al. (2017), an advanced method that accounts for both quantum confinement effects and alloy disorder in determining the effective quantum potential $\frac{1}{u_{e,h}}$. This is achieved by substituting (2) for the conventional Schrödinger equation Filoche et al. (2017). The model integrates in the energy space using the Fermi-Dirac distribution ($f_{n},f_{p}$) and the local density of states for electrons ($N_{c}$) and holes ($N_{v}$) in Eqs. (3) and (4) to compute carrier concentrations, where $E_{fn}$ and $E_{fp}$ denote the quasi-Fermi levels for electrons and holes, respectively.
@@
Figure 3: Simulation models: (a) Compositional map of the combined model incorporating alloy fluctuations and V-defects. Random alloy fluctuations are illustrated in the inset. Indium composition is 17% at V-defect sidewalls QWs, averaged to 29% at c-plane MQWs, and (b) tail states along TD centers and sidewalls, acting as transport and recombination pathways, reducing the turn-on voltage and radiative efficiency, inducing a leakage current below turn-on. (c) Calculated emission spectrum at different current density with the assigned 29% random alloy fluctuation. (d) Band structure showing carrier transport through tail states at low current density, leading to an improved carrier injection at low voltage, which is identified as carrier leakage. (e) Position of tail states in the density of states across the bandgap.
To simulate random alloy fluctuations, we used a random number generator combined with Gaussian weighting methods to distribute indium (In) and gallium (Ga) atoms across the lattice sites within the multiple quantum wells (MQWs). This arrangement effectively simulates local compositional variations and their impact on essential metrics such as bandgap energy and polarization, ensuring accurate device characterization. Our model also accounts for the impact of tail states, which are carrier states below the bandgap that arise from crystal structure imperfections, strain-induced bandgap changes, or impurities. These tail-state effects have been widely discussed in polycrystalline, doped, or amorphous semiconductor materials. Such states provide additional carrier percolation paths, significantly influencing carrier dynamics. The origin of tail states in this paper will be discussed later. To incorporate tail states, we define the density of tail states, $N_{tail}(E)$, represented by a Gaussian distribution [Eq. (12)], as shown in Fig. 3(e). This equation describes the peak density of tail states $N_{tail}$, the energy center $E_{tail}$, and the standard deviation $\sigma_{tail}$.
@@
Here, $N_{(c,v),total}(E)$ replaces $N_{(c,v)}$ in Eqs. (3) and (4), allowing carriers to exist in these tail states and hop through them inside the energy bandgap, as illustrated in Fig. 3(e). Tail states have long been studied in disordered or doped semiconductors. Specifically, tail states have been identified and computed in doped GaAs, where they originate from the local electric fields of ionized impurities Casey Jr and Stern (1976). These tail-state models have been successfully implemented in our software to describe carrier transport in organic and classical semiconductors Kirchartz et al. (2011); Kung et al. (2017); Huang et al. (2020, 2023). Similar functions are now applied here to describe carrier transport in tail states near the TD region.
@@ II.1 Parameter setting
Table 3: The parameters for tail states density and FWHM of Gaussian distribution in threading dislocation lines and V-defect sidewalls. $E_{tail}$ is zero in the setting. The value used here gives a satisfactory fit to the WPE curve shown in Ref. [Zhang et al. (2020)]. The QW region has a random alloy fluctuation, and the $\sigma_{tail}$ changes from 0.4eV to 0.3eV for indium composition varies from 0% to 100%, respectively.
@@
QW region $10^{16}$ 0.4-0.3
Further material parameters, such as bandgap energies, doping concentrations, and recombination coefficients, are provided in Table 1. To account for spatially varying material quality, non-radiative lifetimes ($\tau_{nr}$) were assigned to three distinct Quantum Well (QW) groups: QWs 1-3, QWs 4-6, and QWs 7-9. This physically motivated simplification was adopted because: (1) QWs grown at different stages may exhibit varying defect densities; (2) carrier recombination predominantly occurs in QWs near the p-side (QWs 7-9), making their ($\tau_{nr}$) most critical; and (3) this approach avoids an excessive number of free parameters and overfitting, unlike assigning individual lifetimes to all nine QWs, while still capturing essential spatial variations in non-radiative recombination. A uniform ($\tau_{nr}$) model was found insufficient for accurate IQE fitting. The model’s sensitivity to these grouped lifetimes is further detailed in the Supplementary Material (Fig. S1-S4). The Auger-Meitner coefficient $C_{0}$ is chosen to be significantly larger in the InGaN QW region, as it is well known that alloy disorder substantially increases this coefficient compared to pure GaN Shahmohammadi et al. (2017); Ryu et al. (2020). This choice of $C_{0}$ also provides better agreement with the experimental results discussed later. To accurately model the experimental emission spectrum (peaking around 590-625 nm depending on current density, Ref. [Zhang et al. (2020)]), careful determination of the effective In concentration in the c-plane QWs is crucial, particularly given the TEM-verified QW thickness of  2.5 nm [Ref. Zhang et al. (2020)]. While Ref. [Zhang et al. (2020)] mentions a nominal  40% Indium target, such values are often initial estimates based on Vegard’s law for uniform bulk InGaN bandgaps, which do not fully account for Quantum Confined Stark Effect (QCSE) and random alloy fluctuation in thin QWs. Achieving precise high-In incorporation and sharp interfaces in InGaN QWs is inherently complex, often leading to deviations between nominal targets and actual compositions, as documented in studies on In incorporation challenges and advanced characterization (e.g., Refs. [ Duboz et al. (2023); Liu et al. (2014)]). Furthermore, the presence of random alloy fluctuations means that light emission is often dominated by localized In-rich sites as shown in Fig. 3 (a)). This localization effect contributes to the observed difference between a nominal In target and the effective average composition needed to model the overall spectrum.
Given these constraints, we systematically investigated different initial local In contents within the 2.5 nm thick c-plane QWs to best match the experimental spectrum. The Schrödinger solver is used to calculate the eigenstates and the emission spectrum shown in Fig. 3(c) with the fluctuated potential in the c-plane QW. Notably, the bowing coefficient for the InGaN alloy bandgap is taken as -2 eV [Tsai et al. (2023)]. Our model begins by randomly assigning Indium and Gallium atoms within the QW lattice sites with an initial target local average Indium content (such as 37% in the final result). To determine the actual local indium composition a Gaussian weighting method is employed [Ref. Li et al. (2017)]. This method calculates the local In concentration by applying Gaussian weights to surrounding In and Ga atoms, thereby reflecting the immediate atomic environment. This process inherently considers the influence of the neighboring GaN Quantum Barriers (QBs) where only Ga atoms (zero In) are present near the interface. The resulting spatially varying local Indium composition, exhibiting significant fluctuations, is depicted in Fig. 3(a) (and conceptually in Supplementary Material, Fig. S5(a)). This provided the optimal fit to the experimental emission spectrum from Ref. [Zhang et al. (2020)] with a peak in-plane average indium composition at 29%. Crucially, our model, which includes these random alloy fluctuations and the strong QCSE, demonstrates that the actual light emission predominantly originates from localized In-rich sites where the Indium content can significantly exceed this 29% average, and indeed locally surpass even a 40% nominal value (as illustrated by the high-Indium regions in Fig. 3(a)). While the effective average peak for emission calculation is 29%, the underlying random distribution means local regions with higher Indium concentrations exist and can influence carrier localization and emission spectrum.
@@
For the V-defect structure. The thin (0.67 nm) V-defect sidewall QWs were designed with an indium composition of 17%, along with 3.62 nm thick sidewall QBs. These variations lead to corresponding changes in the bandgap energies and polarization properties of the sidewall QWs, which can impact carrier dynamics and recombination efficiency. In the TD center, within 30 nm regions, there is a high density of deep-level traps and shallow-level tail states. The deep-level trap state parameters are listed in Table 2, while the tail state parameters are shown in Table 3.
@@
Our findings indicate that tail states significantly influence carrier transport Kung et al. (2017); Huang et al. (2020), particularly when strong polarization-induced fields and large band offsets hinder vertical carrier injection into the QWs, as in red LEDs. This disrupts the ideal carrier flow over barriers, as illustrated by the band structure model in Fig. 3(d). The altered energy landscape redirects carriers, increasing leakage to non-radiative pathways and promoting recombination at NR states near dislocations. Furthermore, the inclusion of tail states (Table 3) modifies the density of states (DOS), as shown in Fig. 3(e). The additional energy levels introduced by tail states increase the probability of carrier trapping, enhancing nonradiative recombination in dislocation regions and reducing radiative efficiency. These effects on the DOS highlight the critical role of tail states in shaping the electronic properties and overall optical performance of the LED.
@@
The specific values for parameters such as non-radiative lifetimes and tail state characteristics were determined by optimizing the fit to experimental J-V and IQE data from Ref. [Zhang et al. (2020)], within physically reasonable constraints. While an exhaustive exploration of all parameter combinations is computationally prohibitive, illustrative sensitivity analyses for key parameters (non-radiative lifetimes, different indium compositions for emission peak verification) are provided in the Supplementary Material (Figs. S1-S6) to support our parameterization and validate the model’s behavior. To ensure full reproducibility, the final simulation input file for our publicly available 2D-DDCC solver and a corresponding execution guide are also included in the Supplementary Material.
@@ II.2 Simulation models
It is important to note that the V-defects incorporated in our simulations (e.g., in the third and fourth models) are not treated as isolated entities. To represent realistic device conditions and investigate density effects, our simulations are performed over a 10 $\mu$m wide domain, typically including multiple V-defects (10 V-defects for baseline results, which corresponds to 108 cm-2 density). The similar approach was made in blue and green LED studiesHo et al. (2022). Accurately resolving the intricate current flow around these multiple defects, alongside random alloy fluctuations, necessitates an extremely fine computational mesh (e.g., dx=0.5nm, dy=0.1nm in active regions), leading to approximately 16.6 million nodes per 2D slice. This results in a substantial computational burden, with each full J-V curve simulation requiring over 80 GB of RAM and 1-5 days of processing time on multi-core CPUs. Despite these demands, this multi-defect approach is crucial for a more accurate assessment of V-defect impact. The influence of varying V-defect densities, achieved by changing the number of explicit V-defects within this domain, is further analyzed later.
It is important to note certain simplifications adopted in our 2D simulation model to maintain computational feasibility while focusing on the key physical phenomena governing the device characteristics under investigation. One such simplification is the omission of the InGaN/GaN superlattice (SL) structure typically present beneath the active region in many experimental LED designs, including that described in Ref. [Zhang et al. (2020)].
Including the full SL stack would lead to a prohibitive increase in computational resources (estimated  5x memory, significantly extended time), given our already demanding 2D model which incorporates random alloy fluctuations, multiple V-defects, and tail states within a large domain. Our primary focus is elucidating the 2D carrier transport within the MQW active region itself, including V-defect assisted injection and leakage mechanisms. While SLs undoubtedly influence strain and crystal quality, thereby impacting Internal Quantum Efficiency (IQE), their direct effect on the fundamental turn-on voltage characteristics can be less significant when they primarily serve as underlying quality-enhancement or strain-management layers. For instance, studies on GaN-based LEDs incorporating InGaN/GaN SLs underlying the active region have shown significant improvements in crystal quality and optical output, while the turn-on voltage from current-voltage characteristics remained largely unaffected by the presence of such SLs Zhao et al. (2020); Tao et al. (2018). Tao et. al. shows that more SL even reduce the voltage due to their effect on the V-defect size or densities. This suggests that while crucial for optimizing overall device efficiency through material improvement, the primary electrical injection barriers are still largely dictated by the active MQW, V-defect, and p-n junction properties. Therefore, omitting the detailed SL structure allows us to dedicate computational resources to the complex 2D phenomena within the active region, which are critical for accurately modeling the J-V characteristics and leakage currents. This modeling choice differs from 1D SL simulations [e.g., Ref. Jiang et al. (2019)] that cannot resolve these critical 2D effects essential for accurate J-V fittingHo et al. (2022); Wu et al. (2012). This omission is an acknowledged limitation regarding the absolute prediction of IQE values that are dependent on SL-induced quality improvements, but the core electrical transport phenomena are believed to be robustly captured.
Figure 5: Computations of current density distributions of (a, b, c) electrons and (d, e, f) holes at applied voltages of V = 2.85V and 2.66V for the third model (compositional fluctuations + V-defects, (a, b) and (d, e), and for the fourth model (addition of tail states (c, f)). V-defects improve carrier flow while tail states increase leakage and nonradiative recombination in the defect region. However, holes are then leaking to lower QWs so that more QWs will emit light. Also, a lower $V_{for}$ is needed to reach a current density of 20 A/cm2 compared to the model without tail states.
@@ III.1 Carrier Injection Mechanism Through V-defects
In Fig. 5(a,d), V-defects and random alloy fluctuations are incorporated into the third model. Since the sidewalls of the QWs are semipolar planes with minimal piezoelectric polarization, the V-defects provide an additional channel for carrier injection. As a result, carrier flow into the c-plane QWs is more efficient through the sidewalls of the V-defects. The reduced barriers, caused by the lower piezoelectric polarization, allow for a more concentrated current density at the V-defect sites, where the voltage at a current density of 20 A/cm2 significantly drops to 2.85V. Additionally, carriers trapped along the TD lines may lead to nonradiative recombination losses. Although the forward voltage decreases significantly, it remains higher than the experimental value of approximately 2.1 V at 1 A/cm2. This discrepancy arises because, although the V-defect $\{10\overline{1}1\}$ semipolar plane ($\sim 62^{\circ}$) exhibits nearly zero polarization at the interface, it typically features a thinner QW with a lower indium composition, similar to that of blue/violet MQWs. The theoretical turn-on voltage of blue LEDs is approximately 2.72V ($\hbar\omega=2.72eV$). Consequently, the sidewall QW aids carrier injection once it becomes conductive. Even if the sidewall QW had a higher indium composition (22%) instead of the 17% used, the voltage would decrease slightly but still would not reach the low turn-on voltage due to the increased band offset in the QW and the higher series resistance resulting from the presence of nine deeper sidewall QWs. Therefore, incorporating a tail state current transport model is necessary.
@@
When the voltage is lowered to 2.66V, fewer carriers are injected in the third model due to the absence of defect-mediated pathways, requiring a higher voltage to achieve comparable injection levels (Fig. 5(b, e)). However, in the fourth model, which includes tail states, the current density reaches 20 A/cm2 at this voltage (Fig. 5(c, f)). These tail states are incorporated into the V-defect region in the fourth model, with $N_{tail}$ being higher at the defect center (the location of the dislocation line, Fig. 3(b)). The higher current densities in the TD area (Figs. 5(c) and 5(f)) indicate that these tail states increase the current, particularly at lower voltages, while also leading to greater nonradiative recombination, as the TD area contains fast nonradiative recombination traps. The tail states also create leakage paths toward nonradiative centers, reducing overall efficiency. However, they facilitate carrier injection into the QWs at low voltages, as some leakage current diffuses into the QWs. Once carriers begin injecting into the c-plane quantum well (QW), they screen the polarization field, particularly at the corners of the c-plane/semipolar plane interface. This screening effect reduces the injection barrier in the c-plane region. Comparing Figs. 5(e) and 5(f) for hole injection at the same voltage, we observe slightly stronger vertical injection in the c-plane or semipolar plane, especially at the QW corners. Additionally, the J-V curves in Figs. 7 and 8 indicate that this change is not limited to leakage current below the turn-on voltage: the entire J-V curve shifts to a lower voltage.
@@
Figure 9: (a) and (b) are the linear and log scale IV curves of different tail state Gaussian width $\sigma_{tail}$ and $N_{tail}$ in QB region in TD center or rest V-defect region as defined Fig. 3(b). The tail state setting in QW is kept the same. The reference cases have no tail state in either the QW or QB region. These curves show that increasing $\sigma_{tail}$ or $N_{tail}$ results in a lower turn-on voltage. (c) and (d) are the IQE and WPE curves versus current density, respectively. The larger $\sigma_{tail}$ or $N_{tail}$ leads to a smaller IQE. However, a smaller voltage helps slightly with WPE.
Figure 10: (a) and (b) I–V curves in linear and log scale, respectively, for varying V-defect densities, showing a lower forward voltage at higher defect densities. (c) and (d) are the dependence of IQE and WPE on current density, respectively. Moderate V-defect density enhances efficiency, whereas excessive density leads to nonradiative losses.
Figure 11: (a) IV curves for different nonradiative lifetime, showing that the forward voltage remains nearly unchanged, indicating minimal impact on carrier injection. The three values on nonradiative lifetime are for QW1-3, QW 4-6, and QW 7-9, respectively. (b) WPE as a function of current density, demonstrates greater sensitivity to lifetime variations. The experimental value (Ref.[s]) is from the Ref.[Zhang et al. (2020)].
As discussed in Fig. 8, suitable parameters are adjusted to fit the experimental result closely, especially in the low current density region with a reasonable voltage and IQE. Discussing the influence of these key parameters on the result is important to understanding physics further. As shown in Fig. 9(a), we change the broadening term of $\sigma_{tail}$ in QB in analyzing the influence of leakage current. The tail state setting in QW is kept the same for comparison. As shown in Fig. 9, the reduction of $\sigma_{tail}$ in the TD and V-defect region affects the distribution of tail states, which makes the leakage smaller in the low current density region. The reduction of $N_{tail}$ gives similar effects. Hence, to fit different results from different growth conditions, we will need to adjust these parameters slightly. Since the TD center is the major nonradiative recombination center, adding a large $\sigma_{tail}$ will provide a path for current injection and lead to a larger nonradiative recombination and smaller IQE as shown in Fig. 9(c). The IQE becomes higher as $\sigma_{tail}$ decreases because of less carrier injection into the TD center. However, if we calculate the WPE related to the voltage, the case without tail states has a smaller WPE due to a larger forward voltage. When $\sigma_{tail}$ in QB region is small, further reducing $N_{tail}$ continues to increase the forward voltage. The influence on IQE is slightly changed since the ratio of carrier injection into different QWs changes.
We further investigate the influence of V-defect density, including the tail states as shown in Fig. 10. We can find that smaller defect density has a larger voltage due to less leakage path from the sidewall of V-defect and tail states. Because the current is crowded near the V-defect, we can see an earlier droop behavior due to the Auger recombination from the current crowding. More V-defects help the current injection and diminish current crowding. The forward voltage is also smaller. However, if the V-defect density is too high, the active c-plane area will become smaller, and there will be more nonradiative recombination centers, which will decrease the IQE again. Hence, the WPE reaches the maximum at V-defect density around 1-2 ×$\times 10^{8}$ $\mathrm{cm}^{-2}$. Similar observation has been reported in the early work for blue and green LEDs Ho et al. (2022)
Finally, the nonradiative lifetime resulting from the material growth condition plays a key role in WPE or IQE, as shown in Fig. 11. A shorter nonradiative lifetime will lead to a smaller IQE but does not affect the I-V curve significantly. The nonradiative lifetime may result from point defect densities. The different designs of super lattices for the pre-strained layer to open the V-defect Jiang et al. (2019) may also lead to different material quality, which all present different lifetime. These will affect the IQE of the device directly.